# Pathology-CoT: Learning Visual Chain-of-Thought Agent from Expert Whole Slide Image Diagnosis Behavior

### v1

@@ 2.3 Pathology-o3 and learned behavior can generalize on external validation dataset
To test for generalization, the Behavior Predictor, trained on Stanford Medicine data, was evaluated on a completely independent external dataset: the lymph node in colon adenocarcinoma (LNCO2) cohort. This challenging cohort consists of 1,245 whole-slide images from Sweden, introducing domain shifts from different scanners and magnifications. Despite these variations, pathology-o3 trained on United State pathologist behavior maintained robust performance, markedly outperforming other leading VLMs on this benchmark (Figure 1a).
Quantitatively, our agent achieved an accuracy of 69.4%, a precision of 62.9%, and a near-perfect recall of 97.6%. These results demonstrate a significant leap in performance; for instance, pathology-o3’s precision was more than double that of the next-best model (Gemini at 23.5%), and its exceptional recall highlights a strong ability to reliably identify positive cases.
@@ 2.4 Pathology-o3 learns where to look
We next examined two key properties of our behavior-guided agent: the robustness of its learned policy and the fundamental importance of navigation itself. We first evaluated robustness by permuting the order of ROI presentation (forward, reverse, and random). The agent’s performance varied only minimally across these settings (Figure 4a, left), indicating that it synthesizes evidence effectively without dependence on a fixed sequence. This suggests that as long as diagnostically relevant regions are identified, the specific viewing habits of different pathologists can be learned and successfully deployed within an agentic system.
Second, we analyzed the agent’s performance when we capped the maximum number of ROIs it could inspect ($k$, analogous to CoT length). As shown in Figure 4a (right), performance improved steadily as $k$ increased, with notable gains in recall while preserving high precision. This finding demonstrates that when provided with a sequence of diagnostically relevant regions, VLMs are highly effective at synthesizing the evidence to reach an accurate conclusion.
This proficiency in analysis raises a critical question: if VLMs are strong analyzers, what limits their performance in WSI-level tasks? To investigate, we evaluated their out-of-the-box navigation capabilities. Rather than being guided by our Behavior Predictor, we let VLMs to select their own ROIs. We then compared their navigation against a senior attending’s path using two metrics (described in Methods): completeness, which represents the fraction of expert-viewed regions recovered; and efficiency, which represents the proportion of selected regions that were clinically relevant. As shown in Figure 4b (upper), unguided VLMs performed poorly on both measures. Qualitatively, their ROI selections were scattered and low-yield, whereas our behavior-guided agent concentrated on clinically salient sites that closely overlapped with the expert’s annotations (Figure 4b, lower). These results reveal a fundamental gap between analysis and navigation. The strong analytical ability of VLMs reflects training on large corpora of static, pre-cropped images, whereas their failure to navigate stems from the absence of data capturing continuous, visually guided interaction with complex environments. This finding underscores our data-centric approach: building effective agents requires digitizing the tacit, procedural knowledge of expert navigation into a structured format suitable for learning.
Figure 3: Quantitative and qualitative results on the external LNCO2 validation cohort. a Bar charts comparing Accuracy, Precision, and Recall of Pathology-CoT with other VLM backbones. b Qualitative examples of the model’s output on two slide images, with highlighted regions and corresponding textual descriptions.
@@ 2.5 Practicality, scalability, and the value of learned behavior
Finally, we analyzed the practical viability and scalability of our agentic framework. Figure 5a provides a cost and latency snapshot for a representative WSI analysis, comparing a large, powerful model Gemini-2.5-pro with a smaller, faster one Gemini-2.5-flash. The results demonstrate that the agentic workflow is already feasible with current backbones. More importantly, our framework is modular by design, allowing the core reasoning engine (the VLM) to be upgraded as foundation models evolve. This design ensures the system is future-proof: it can readily incorporate next-generation models to become progressively faster, cheaper, and more capable over time.
Next, we investigated the value of the behavior-aligned data itself. We compared three conditions across multiple VLM backbones (Figure 5b): (i) Non-agent is a baseline without guidance and diagnosis with only look at the image once; (ii) Agent guided by real behavior from a senior attending GI pathologist (a practical upper bound) as we shown in . and (iii) Agent guided by learned behavior from the behavior bank; The addition of behavioral guidance yielded a profound improvement in performance. Compared to the baseline, guidance from the learned viewing policy increased precision by 11.8% and recall by 17.9%. The gains from real expert behavior were similarly large, boosting precision by 11.9% and recall by 19.1%. This translated to a marked increase in overall accuracy, improving it by 9.2% and 7.1%, respectively. Critically, our agent’s learned behavior closes most of the performance gap to real behavior. This has a powerful implication for scalability: instead of requiring continuous, large-scale expert data collection, we can train a compact and efficient Behavior Predictor on a limited dataset. This predictor effectively distills the expert’s viewing policy, creating a scalable asset that can generate human-aligned guidance for any VLM on any number of slides, reinforcing our central thesis that the most effective path to a capable pathology agent is by first building the right data engine.
Figure 5: Comparison of different ‘thinking’ modes and cost-benefit analysis. a A table comparing cost, time, and performance between high-cost reasoning model Gemini-2.5-pro and low cost. b Performance comparison of ‘Non-agent’, ‘Agent guided by real behavior’, and ‘Agent guided by learned behavior’ modes across various VLMs.
@@ 3 Discussion
The rapid maturation of vision–language models is creating a new frontier for artificial intelligence in medicine, shifting the paradigm from task-specific classifiers to versatile, reasoning-based agents. Digital pathology, with its gigapixel canvases and complex, multi-step diagnostic workflows, is a particularly fitting domain for this agentic approach. However, the primary obstacle to building capable pathology agents has since been bottlenecked by high quality data annation rather than modeling: supervision that is not only large-scale but also deeply aligned with the procedural, tacit knowledge of clinical experts.
@@
The central challenge is that this expert knowledge is not readily available in a structured format. While modern pathology viewers generate vast digital logs of user interactions, this raw behavioral data is high-frequency, noisy, and enormously long. It captures the outcome of an expert’s thought process but not the discrete, intention-driven steps that are amenable to model training. This results in a fundamental "analysis–navigation gap": while VLMs are trained on vast corpora of static images and are thus adept at analyzing pre-selected regions, they are not trained on the dynamic, interactive data needed to learn the procedural skill of navigating a slide to find those regions in the first place.
@@
To bridge this gap, we developed the AI Session Recorder, a data engine that digitizes the tacit knowledge of expert navigation into our Pathology-CoT dataset. We demonstrated the value of this behavior-aligned data by building a simple agent, Pathologist-o3, as a controlled testbed. Our results show that this data provides a universal, model-agnostic performance multiplier, significantly improving the diagnostic accuracy of a wide range of VLM backbones. We found that the learned viewing policy is robust, generalizing effectively to an entirely different hospital system with different scanners and patient populations. Further analysis revealed that this policy is not only robust to variations in viewing habits but is also essential for effective diagnosis, as leading VLMs without this guidance fail at the fundamental task of navigation. Finally, we established that this data-centric approach is both practical and scalable; the expert viewing policy can be distilled into a compact, efficient model that provides nearly all the benefits of real expert guidance, creating a powerful asset for building the next generation of capable, human-aligned pathology agents.
@@
The study has several limitations, primarily centered on our data-centric methodology. First, our findings on viewing policy, such as robustness to viewing order, are specific to the "search-and-find" task of metastasis detection. This may not generalize to tasks requiring a systematic, sequential evaluation, like architectural pattern grading in prostate cancer or counting mitoses along a tumor boundary, where the viewing path itself is a critical component of the diagnosis. To begin addressing this, we are continuously expanding our public dataset with behavior logs from a wider range of pathological tasks, with ongoing data releases available on our project website. Second, our efficient human-in-the-loop workflow for generating reasoning labels introduces a potential for AI anchoring bias, where the initial AI-generated text could subtly influence the expert’s final revision. Finally, a practical challenge lies in the data source itself. Unlike standardized web-scale corpora, pathologist viewing logs are heterogeneous, with different software platforms using varied and often unstructured formats. Tapping into this data requires an initial engineering effort for each system. However, the significant advantage of our approach is that once this integration is complete, data collection becomes entirely passive, imposing no additional burden on the pathologists’ routine clinical workflow.
@@ 4 Method
Figure 6: Pathology-CoT Dataset Creation Pipeline. Left: Pathologist interaction data (viewports, navigation actions like zoom-in/panning, task context) is collected. Center: Generating the ‘Thinking Label’. Contextual information (task, current view, next action) is used to prompt a large language model (LLM, e.g., Gemini) to generate candidate reasoning. A pathologist then reviews, corrects, or approves this output to create the final ‘Thinking Label’. Right: The collected actions and generated thinking labels are merged with corresponding image patches (thumbnails, zoomed regions) and system prompts into a structured conversational format suitable for model training or in-context learning.
@@ 4.1 Data Collection: Pathologist Behavior
We collected interaction data from eight pathologists with varying experience levels (four attendings, two fellows, and two residents) analyzing 137 diagnostic Whole Slide Images (WSIs) from 25 cases of colorectal cancer (CRC) lymph node (LN) metastasis. The data collection setup, using the nuclei.io software, recorded the following:
@@
This raw data captures the dynamic, multi-modal behavior of expert pathologists during routine diagnostic tasks (Figure 6, left panel).
Figure 7: Sankey Diagram of Diagnostic Attention Flow. The diagram illustrates how two primary examination methods, Inspect (broad-field scan) and Peek (high-magnification look), are applied to analyze various histological and cytological features during lymph node metastasis diagnosis. The flow size represents the relative proportion of diagnostic attention directed from a specific method to each feature.
@@ 4.2 AI Session Recorder: Discretizing Behavior to Extract ROIs
The foundation of our approach is the AI Session Recorder, a computational pipeline designed to transform the raw interaction data gathered from pathologists—a high-frequency, continuous, and inherently noisy stream of timestamped viewports (Figure 7a) into structured, agent-ready data. Pathologists’ navigation is rapid and complex; a single slide analysis can generate an average of 257.3 viewport events. Directly using this raw data for model training would be chaotic and impractical. If each of these 257 viewports were treated as a separate 1024x1024 patch, it would generate over 500K tokens from the visual modality alone. This volume is prohibitive for prominent large models, and could severely degrade performance.
The AI Session Recorder processes this raw stream using a series of heuristics inspired by the physical use of a microscope, which has discrete, not continuous, magnification levels. We analyze the temporal patterns of viewport changes, as shown in Figure 7c, to segment the continuous stream into discrete actions. We define two primary types:
@@
These initial actions are then filtered to remove low-magnification overviews, merged if they have high spatial overlap (IoU > 0.8), and pruned to prioritize the most specific, high-magnification views. Finally, each resulting action’s magnification level is binned to the nearest standard microscope objective, and its bounding box is normalized to a standard size. For instance, a 12x view becomes a <10x-inspect> action. As shown in Figure 7(b), this converts the raw data into a clean, compact sequence of behavioral commands. This abstracted action sequence, composed of a command type and a standardized ROI box, serves as the basis for both training our agent and generating reasoning labels.
@@ 4.3 AI Session Recorder: Rationale Generation
    Pathologist Review and Correction: Critically, the AI-generated draft is presented to a human pathologist in a lightweight interface for rapid validation, as shown in Supplementary Figure S3. The interface displays the relevant thumbnail, the zoomed-in ROI, and an optional cytology-level crop alongside the generated text for "Thumbnail Impression," "Why Zoom," and "Findings." The pathologist can then efficiently accept, edit, or reject the text to ensure its clinical accuracy and alignment with their actual reasoning process.
@@ 4.4 The Pathology-CoT Dataset
The curated rationale and impression components are structured alongside the corresponding state information to form the final Pathology-CoT dataset. As illustrated in Figure 6 (right panel), the data can be formatted into sequences of state-action tuples $(s_{t},a_{t+1}^{\text{GT}})$, or arranged in a conversational format suitable for prompting VLMs using In-Context Learning.
The final output of our pipeline is the Pathology-CoT dataset, a structured collection of expert-validated, agent-ready training data. The dataset is derived from 10.6 hours of recorded interactions from eight pathologists at Stanford Hospital (including attendings, fellows, and residents) analyzing WSIs for the task of colorectal cancer lymph node metastasis. The current version contains 921 sessions and 5,222 conversational rounds. Each round consists of a behavioral command, a standardized ROI, and the corresponding expert-validated reasoning. The reasoning is richly detailed, with broad <inspect> actions accompanied by an average of 152 words of description, and high-magnification <peek> actions averaging 82 words on cellular-level findings.
4.5 Pathologist-o3 Agent Framework and Evaluation

### v2

@@ 2.3 Pathology-o3 and learned behavior can generalize on external validation dataset
To test for generalization, Pathology-o3 and the Behavior Predictor, trained on Stanford Medicine data, was evaluated on a completely independent external colon adenocarcinoma lymph node metastasis dataset: LNCO2 [27]. This challenging cohort consists of 321 whole-slide images from Sweden, introducing domain shifts from different scanners and magnifications. Despite these variations, Pathology-o3 trained on United State pathologist behavior maintained robust performance, substantially outperforming other leading VLMs on this benchmark (Figure 3a).
Quantitatively, our Pathology-o3 agent AI achieved an accuracy of 69.4%, a precision of 62.9%, and a near-perfect recall of 97.6%. These results demonstrate a significant leap in performance. For instance, Pathology-o3’s precision was more than double that of the next-best model (Gemini at 23.5%), and its exceptional recall highlights a strong ability to reliably identify positive cases. More detail on performance metrics are shown in Table S3.
@@
Figure 3: Quantitative and qualitative results on the external LNCO2 validation cohort. a Bar charts comparing Accuracy, Precision, and Recall of Pathology-CoT with other VLM backbones. Error bars represent 95% bootstrap confidence intervals. b Qualitative examples of the model’s output on two slide images, with highlighted regions and corresponding textual descriptions.
@@ 2.4 Pathology-o3 learns where to look
We next examined two key properties of our behavior-guided agent: the robustness of its learned policy and the fundamental importance of navigation itself. We first evaluated robustness by permuting the order of ROI presentation (forward, reverse, and random). The agent’s performance varied only minimally across these settings (Figure 4a, left), indicating that it synthesizes evidence effectively without dependence on a fixed sequence. This suggests that when identify lymph node metastasis, as long as diagnostically relevant regions are identified, the different viewing habits of pathologists can be learned and successfully deployed within an agentic system.
Second, we analyzed the agent’s performance when we capped the maximum number of ROIs it could inspect, analogous to CoT length. As shown in Figure 4a (right), performance improved steadily as thinking process getting longer, with notable gains in recall while preserving high precision. This finding demonstrates that when provided with a sequence of diagnostically relevant regions, VLMs are highly effective at synthesizing the evidence to reach an accurate conclusion.
This proficiency in synthesizing evidence, however, is entirely contingent upon the VLM being provided with diagnostically relevant regions. This allowed us to directly test our central hypothesis: that the ability to navigate and locate these regions is a distinct, learned skill derived from expert viewing behavior—a skill for which current VLMs lack the necessary training data. To investigate this, we evaluated VLM out-of-the-box navigation capabilities. Rather than being guided by our Behavior Predictor, we let VLMs to select their own ROIs. We then compared their navigation against a senior attending pathologist’s path using two metrics (described in Methods): completeness, which represents the fraction of expert-viewed regions recovered; and efficiency, which represents the proportion of selected regions that were clinically relevant. As shown in Figure 4b (upper), unguided VLMs performed poorly on both measures. Qualitatively, their ROI selections were scattered and low-yield, whereas Pathology-o3 concentrated on clinically salient sites that closely overlapped with the an attending pathologist’s real behavior (Figure 4b, lower). These results reveal a fundamental gap between analysis and navigation. The strong analytical ability of VLMs reflects training on large corpora of static, pre-cropped images, whereas their failure to navigate stems from the absence of data capturing continuous, visually guided interaction with complex environments. This finding reinforces our data-centric approach: building effective pathology agents requires digitizing the tacit, procedural knowledge of expert navigation into a structured format suitable for learning.
@@ 2.5 Practicality, scalability, and the value of learned behavior
Finally, we analyzed the practical viability and scalability of Pathology-o3. Figure 5a provides a cost and latency snapshot for a representative WSI analysis, comparing a large, powerful model Gemini-2.5-pro with a smaller, faster one Gemini-2.5-flash. The results demonstrate that Pathology-o3 is already feasible with current backbones. More importantly, our framework is modular by design, allowing the core reasoning engine (the VLM) to be upgraded as lastest foundation models evolve. This design ensures the system is future-proof: it can readily incorporate next-generation models to become progressively faster, cheaper, and more capable over time.
Next, we investigated the value of the behavior-aligned data itself. We compared three conditions across multiple VLM backbones (Figure 5b): (i) Non-agent is a baseline without guidance and diagnosis with only look at the WSI thumbnail once (similar to PathChat+ [26] approach); (ii) Pathology-o3 guided by real behavior from a senior attending GI pathologist (a practical upper bound) as we shown in Figure 4b; and (iii) Pathology-o3 guided by learned behavior from the behavior predictor. The addition of behavioral guidance yielded a profound improvement in performance. Compared to the baseline, guidance from the learned viewing policy increased precision by 11.8% and recall by 17.9%. The gains from real expert behavior were similarly large, boosting precision by 11.9% and recall by 19.1%. This translated to a marked increase in overall accuracy, improving it by 9.2% and 7.1%, respectively. Critically, Pathology-o3’s learned behavior closes most of the performance gap to real behavior. This has a powerful implication for scalability: instead of requiring continuous, large-scale expert data collection, we can train a compact and efficient behavior predictor on a limited dataset. This predictor effectively distills the expert’s viewing policy, creating a scalable asset that can generate human-aligned guidance for any VLM, reinforcing our central hypothesis that the most effective path to a capable pathology agent is by first building the right data engine.
Figure 5: Comparison of different “thinking” modes and cost-benefit analysis. a A table comparing cost, time, and performance between high-cost reasoning model Gemini-2.5-pro and low cost. b Performance comparison of “Non-agent”, “Agent guided by real behavior”, and “Agent guided by learned behavior” modes across various VLMs.
@@ 3 Discussion
The rapid maturation of vision–language models is creating a new frontier for artificial intelligence in medicine, shifting the paradigm from task-specific classifiers to versatile, reasoning-based agents. Digital pathology, characterized by gigapixel images and multi-step diagnostic workflows, is especially suited to an agentic approach that can automate diagnostic reasoning and accelerate clinical and scientific discovery. However, the primary obstacle to developing capable pathology agents has shifted from modeling limitations to the availability of high-quality data—specifically, supervision that is not only large-scale but also deeply aligned with the procedural and tacit knowledge of clinical experts.
@@
The central challenge is that this expert knowledge is not readily available in a structured format. While modern pathology viewers generate vast digital logs of user interactions, this raw behavioral data is high-frequency, noisy, and enormously long. It captures the outcome of an expert’s thought process but not the discrete, intention-driven steps that are amenable to model training. This results in a fundamental “analysis–navigation gap”: while VLMs are trained on vast corpora of static images and are thus adept at analyzing pre-selected regions, they are not trained on the dynamic, interactive data needed to learn the procedural skill of navigating a slide to find those regions in the first place.
@@
To bridge this gap, we developed the AI Session Recorder, a data engine that digitizes the tacit knowledge of expert navigation into our Pathology-CoT dataset. We demonstrated the value of this behavior-aligned data by building a simple agent AI, Pathology-o3, as a controlled testbed. Our results show that this data provides a universal, model-agnostic performance multiplier, significantly improving the diagnostic accuracy of a wide range of VLM backbones. We found that the learned viewing policy is robust, generalizing effectively to an entirely different hospital system with different scanners and patient populations. Further analysis revealed that this policy is not only robust to variations in viewing habits but is also essential for effective diagnosis, as leading VLMs without this guidance fail at the fundamental task of navigation. Finally, we concluded that this data-centric approach is both practical and scalable; the expert viewing policy can be distilled into a compact, efficient model that provides nearly all the benefits of real expert guidance, creating a powerful asset for building the next generation of capable, human-aligned pathology agents.
Our approach is fundamentally different from other pathology AI agent methods like PathChat+ [26], which delegates the selection of where to zoom and why to an VLM “supervisor” such as GPT-4o. This strategy is limited because the procedural viewing policy of a pathologist is tacit, experience-based knowledge that is not documented in textbooks or on the internet, making it difficult for VLMs to acquire. Consequently, the navigation paths generated by the VLM supervisor often diverge from expert trajectories and lack diagnostic utility, as we illustrate in Sec 2.4. In contrast, by directly learning from recorded expert navigation and pairing it with pathologist-verified rationales, our method provides explicit supervision for both the “where” and the “why,” enabling truly human-aligned behavior and stronger performance.
@@
While this study introduced the powerful “Pathology-CoT” framework, this work still has several limitations, primarily centered on our data-centric methodology. First, our findings on viewing policy, such as robustness to viewing order, are specific to the "search-and-find" task of metastasis detection. This may not generalize to tasks requiring a systematic, sequential evaluation, like architectural pattern grading in prostate cancer or counting mitoses along a tumor boundary, where the viewing path itself is a critical component of the diagnosis. To begin addressing this, we are continuously expanding our public dataset with behavior logs from a wider range of pathological tasks, with ongoing data releases available on our project website. Second, our efficient human-in-the-loop workflow for generating reasoning labels introduces a potential for AI anchoring bias, where the initial AI-generated text could subtly influence the expert’s final revision. Finally, a practical challenge lies in the data source itself. Unlike standardized web-scale corpora, pathologist viewing logs are heterogeneous, with different software platforms using varied and often unstructured formats. Tapping into this data requires an initial engineering effort for each system. However, the significant advantage of our approach is that once this integration is complete, data collection becomes entirely passive, imposing no additional burden on the pathologists’ routine clinical workflow.
@@ 4.1 Data Collection: Pathologist Behavior
This study was approved by the Institutional Review Board (IRB #857258) of the Perelman School of Medicine at the University of Pennsylvania. We collected interaction data from eight pathologists with varying experience levels (four attendings, two fellows, and two residents) analyzing 137 diagnostic Whole Slide Images (WSIs) from 25 cases of colorectal cancer (CRC) lymph node (LN) metastasis. The data collection setup, using the nuclei.io software, recorded the following:
@@
This raw data captures the dynamic, multi-modal behavior of expert pathologists during routine diagnostic tasks (Figure S4, left panel).
@@ 4.2 AI Session Recorder: Discretizing Behavior to Extract ROIs
The foundation of our approach is the AI Session Recorder, a computational pipeline designed to transform the raw interaction data gathered from pathologists—a high-frequency, continuous, and inherently noisy stream of timestamped viewports (Figure S1) into structured, agent-ready data. Pathologists’ navigation is rapid and complex; a single slide analysis can generate an average of 257.3 viewport events. Directly using this raw data for model training would be chaotic and impractical. If each of these 257 viewports were treated as a separate 1024x1024 patch, it would generate over 500K tokens from the visual modality alone. This volume is prohibitive for prominent large models, and could severely degrade performance.
The AI Session Recorder processes this raw stream using a series of heuristics inspired by the physical use of a microscope, which has discrete, not continuous, magnification levels. We analyze the temporal patterns of viewport changes, as shown in Figure S1, to segment the continuous stream into discrete actions. We define two primary types:
@@
These initial actions are then filtered to remove low-magnification overviews, merged if they have high spatial overlap (IoU > 0.8), and pruned to prioritize the most specific, high-magnification views. Finally, each resulting action’s magnification level is binned to the nearest standard microscope objective, and its bounding box is normalized to a standard size. For instance, a 12x view becomes a <10x-inspect> action. As shown in Figure S1, this converts the raw data into a clean, compact sequence of behavioral commands. This abstracted action sequence, composed of a command type and a standardized ROI box, serves as the basis for both training our agent and generating reasoning labels.
@@ 4.3 AI Session Recorder: Rationale Generation
    Pathologist Review and Correction: Critically, the AI-generated draft is presented to a human pathologist in a lightweight interface for rapid validation, as shown in Figure S3. The interface displays the relevant thumbnail, the zoomed-in ROI, and an optional cytology-level crop alongside the generated text for "Thumbnail Impression," "Why Zoom," and "Findings." The pathologist can then efficiently accept, edit, or reject the text to ensure its clinical accuracy and alignment with their actual reasoning process.
@@ 4.4 The Pathology-CoT Dataset
The curated rationale and impression components are structured alongside the corresponding state information to form the final Pathology-CoT dataset. As illustrated in (Figure S4, right), the data can be formatted into sequences of state-action tuples, or arranged in a conversational format suitable for prompting VLMs using In-Context Learning.
The final output of our pipeline is the Pathology-CoT dataset, a structured collection of expert-validated, agent-ready training data. The dataset is derived from 10.6 hours of recorded interactions from eight pathologists at Stanford Hospital (including attendings, fellows, and residents) analyzing WSIs for the task of colorectal cancer lymph node metastasis. The current version contains 921 sessions and 5,222 conversational rounds. Each round consists of a behavioral command, a standardized ROI, and the corresponding expert-validated reasoning. The reasoning is richly detailed, with broad <inspect> actions accompanied by an average of 152 words of description, and high-magnification <peek> actions averaging 82 words on cellular-level findings. We also show the detailed distribution of the dataset using sankey plot as shown in Figure S5.
4.5 Pathology-o3 Agent Framework and Evaluation